OpenStudy (anonymous):

Use properties of rational numbers to simplify the expression. 132.06 + 0.12 + 30 135.18 162.18 163.26 162.12

OpenStudy (anonymous):

Who are you?

OpenStudy (anonymous):

igreen2

OpenStudy (anonymous):

.-. Add 12 to the '6' in the decimal, and add 3 to the tens place.

OpenStudy (unklerhaukus):

132.06 + 0.12 + 30 = (132 + 0.06) + 0.12 + 30 = (132 + 30) + (0.06 + 0.12) =

OpenStudy (anonymous):

0.06 + 0.12 = ? 132 + 30 = ?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

Sup Jay.

OpenStudy (unklerhaukus):

132.06 + 0.12 + 30 = (132 + 0.06) + 0.12 + 30 = (132 + 30) + (0.06 + 0.12) = 100 + (32 + 30) + 0.01(6+12) =

OpenStudy (anonymous):

6 + 12 = ? @iGreen2

OpenStudy (anonymous):

18

OpenStudy (anonymous):

Yes, so 0.06 + 0.12 = 0.18

OpenStudy (anonymous):

Now, can you add 132 + 30?

OpenStudy (anonymous):

162

OpenStudy (anonymous):

Yes, so what's our final answer?

OpenStudy (anonymous):

thnx

OpenStudy (unklerhaukus):

are you the same person?

OpenStudy (anonymous):

As iGreen? @UnkleRhaukus

OpenStudy (anonymous):

And who are you talking to?

OpenStudy (anonymous):

yes i am 2 accounts

OpenStudy (anonymous):

just made it

OpenStudy (anonymous):

Oh, what's your other account?

OpenStudy (anonymous):

no this is it

OpenStudy (unklerhaukus):

gaming the system is abuse

OpenStudy (anonymous):

Who's gaming the system?

OpenStudy (unklerhaukus):

any one person with multiple accounts

OpenStudy (anonymous):

No, that's not gaming the system..unless you're medal spamming your other account and stuff like that.